OpenStudy (j2lie):

How to find the vertex, domain, and range of y=-2x^2+8x-3?

OpenStudy (j2lie):

@agent0smith

OpenStudy (j2lie):

the domain is all real numbers, because you can use any number.

OpenStudy (agent0smith):

\[\large x = -\frac{ b }{ 2a }\]to find the x coordinate of the vertex

OpenStudy (j2lie):

What's an x coordinate.

OpenStudy (j2lie):

x will be 4?

OpenStudy (agent0smith):

An x coordinate is... an x coordinate. yes, the x coordinate of the vertex is 4. Plug that into the equation for y, to find the y coordinate.

OpenStudy (j2lie):

it's 5.

OpenStudy (agent0smith):

-2(4)^2 + 8(4) - 3 = ?

OpenStudy (j2lie):

no i did it like this f(2)=-2(2)^2+8(2)-3.

OpenStudy (agent0smith):

why?

OpenStudy (j2lie):

because I checked online and it told me to do that.

OpenStudy (agent0smith):

What told you to do that? We already found x=4. The x coordinate of the vertex is 4. Plug x=4 into the equation for y, to find the y coordinate. You plugged in x=2...

OpenStudy (j2lie):

-2 times 2 equal =-4.

OpenStudy (agent0smith):

Plug in x=4: y = -2(4)^2 + 8(4) - 3 = ?

OpenStudy (j2lie):

y = 21?

OpenStudy (agent0smith):

Oh crap, i just realized that x = -b/2a = -8/(2*-2) = 2, somehow didn't notice that earlier You're right that x=2, y=5, so the vertex is (2, 5)

OpenStudy (j2lie):

ha I told you I'm better than you.

OpenStudy (agent0smith):

lol okay

OpenStudy (j2lie):

I'm just kidding.